40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 15 12:28:03 2016, gautam, Update, General, Something has gone wrong - was there a power outage? Vac_15Feb2016.png
    Reply  Mon Feb 15 13:09:33 2016, Koji, Update, General, Something has gone wrong - was there a power outage? 
       Reply  Tue Feb 16 09:13:39 2016, Steve, Update, General, there was a power outage & EQ 20160215_power_outage.png
          Reply  Tue Feb 16 15:02:19 2016, ericq, Update, General, Power Glitch recovery 
             Reply  Tue Feb 16 23:42:22 2016, gautam, Update, General, Power Glitch recovery - arms recovered 
                Reply  Wed Feb 17 13:45:15 2016, ericq, Update, General, No monit on frontends 
             Reply  Thu Feb 25 11:02:36 2016, Steve, Update, General, Power Glitch again powerGlitch.png
                Reply  Thu Feb 25 11:32:04 2016, gautam, Update, General, Power Glitch again 
          Reply  Mon Feb 22 08:45:46 2016, Steve, Update, SUS, EQ 4.3 grabs ITMX-UL magnet EQ4.3LucerneV.pngITMX-UL.png
             Reply  Mon Feb 22 23:00:21 2016, ericq, Update, SUS, ITMX free, green is flashing 
                Reply  Tue Feb 23 09:51:34 2016, Steve, Update, SUS, ITMX free ITMX_FREED.pngITMX_FREE.png
                   Reply  Tue Feb 23 17:46:09 2016, gautam, Update, SUS, EQ  SUS_Summary_23Feb.png
             Reply  Tue Apr 5 08:55:51 2016, Steve, Update, SUS, local EQ 3.1m eq3.1Valencia.png
                Reply  Wed May 25 08:32:55 2016, Steve, Update, SUS, local EQ 3.5m 3.5Cucam.pnglocal3.5cucam.png
                   Reply  Fri Jun 3 14:24:13 2016, Steve, Update, SUS, local EQ 3.1 m eq3.1mMuscoyCa.png
                      Reply  Fri Jun 10 09:52:51 2016, Steve, Update, SUS, EQ 5.2 mag Borego Spring 5.2mBoregoSpring.pngeqBSs.png
                         Reply  Mon Jun 20 09:36:10 2016, Steve, Update, SUS, EQ 3.1 m Big Bear Lake, CA eq3.1BigBearLakeCA.png
Message ID: 11990     Entry time: Mon Feb 15 12:28:03 2016     Reply to this: 11991
Author: gautam 
Type: Update 
Category: General 
Subject: Something has gone wrong - was there a power outage? 

I came into the 40m a few minutes ago, and noticed the following (approximately in this order):

  • The striptool plots projected onto the wall were gone, even though the projector seemed to be working fine
  • There was no light at all in the IFO 
  • There was an incessnt beeping noise coming from inside the lab.

To investigate further, I checked today's summary pages, and whatever caused this, happened around 730am today morning (approx 5 hours ago). I also saw that all the watchdogs were tripped, except MC3, BS and SRM. 

I then tracked down the beeping - I believe that it is coming from Megatron.(in fact, it is coming from the Jetstor..) 

I also found that the PSL is OFF, and the Marconi, though ON, has the display parameters set to values that I normally see when it is first turned ON (i.e. the carrier frequency is 1200MHz, the output is -140dBm etc - this is what led me to suspect that somehow the power connection was interrupted? As far as the workstation computers are concerned, I don't think ROSSA was affected, but pianosa is frozen and donatella is at the login screen. The CDS overview MEDM screen refuses to load correctly (though some of the other MEDM screens are working fine). I'm not entirely sure how to go about fixing all of this, so for now, I'm leaving the PSL off and I've shutdown the remaining watchdogs.

It just occurred to me to check the status of the vacuum - the MEDM screen seems to suggest everything is fine (see Attachment #1). I went down to the X end to do a quick check on the status of the turbo pumps and everything looks normal there...

Attachment 1: Vac_15Feb2016.png  33 kB  | Hide | Hide all
Vac_15Feb2016.png
ELOG V3.1.3-